# Expanding Chatbot Knowledge in Customer Service: Context-Aware Similar Question Generation Using Large Language Models

### v2

@@ E.4.1 Approximation Guarantee
E.4.2 Complexity Analysis
  • Per Iteration: Computing the marginal gain for all $n$ candidates requires $O(nk)$, where $k$ is the size of the selected subset $S$.
  • Number of Iterations: The algorithm runs for at most $O(n)$ iterations, as each iteration selects one element.
  • Overall Complexity: The total complexity is $O(n^{2}k)$. With precomputed distances, the complexity reduces to $O(n^{2})$ at the cost of $O(n^{2})$ storage.
E.4.3 Practical Implications
The greedy algorithm balances solution quality and efficiency, providing near-optimal results with manageable computational overhead for moderate-sized datasets.

### v3

@@ D.4.1 Approximation Guarantee
D.4.2 Complexity Analysis
The computational complexity of the greedy algorithm is analyzed as follows. Computing the marginal gain for all $n$ candidates in each iteration requires $O(nk)$ operations, where $k$ represents the size of the selected subset $S$. The algorithm executes at most $O(n)$ iterations, as each iteration selects one element. Consequently, the total complexity is $O(n^{2}k)$. With precomputed distances, the complexity reduces to $O(n^{2})$ at the cost of $O(n^{2})$ storage. The greedy algorithm balances solution quality and efficiency, providing near-optimal results with manageable computational overhead for moderate-sized datasets.